The point on the curve y square is equal to X tangent makes 45 degree angle with x axis is

Grade:12

1 Answers

Arun
6 years ago
Dear Vinay
y^2 = x
hence
on differentiating
2y dy / dx = 1
dy/dx = 1/2y
now 1/2y = tn 45 = 1
y = ½
hence
(½)^2 = x
x = ¼
hence point = (1/4 , 1/2)
 
 
Regards
Arun (askIITians forum expert)